Determine the binding energy of an O-16 nucleus. The O-16 nucleus has a mass of 15.9905 amu. A proton has a mass of 1.00728 amu, a neutron has a mass of 1.008665 amu, and 1 amu is equivalent to 931 MeV of energy.
Answer #1
An oxygen-16 nucleus has 8 protons and 8 neutrons. The sum of the masses of these particles, if they were isolated (not bound together in a nucleus) would be:

8*(1.00728 AMU) + 8*(1.0088665AMU) = 16.12756 AMU.

The observed mass of an O-16 nucleus is 15.9905 AMU, the difference between this value, and the sum of the masses of the isolated nucleons is called the "mass defect", and is equal to the binding energy of the nucleus.

BE = (16.12756 - 15.9905)AMU = 0.13706 AMU.

Using the conversion you were given,

0.13706 AMU * 931MeV/AMU = 127.6 MeV

1 MeV = 1.602 * 10^-13 J, so in joules, the binding energy is 2.044*10^-11 J

(Depending on the convention used, the binding energy is sometimes expressed as a negative number. The energy of the O-16 nucleus is less than that of the isolated protons and neutrons, so it is more stable than a collection of isolated particles.)